OpenStudy (anonymous):

2(x+8)=4x-2 is it -6

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

2x+16=4x-2

OpenStudy (anonymous):

i got 3

OpenStudy (anonymous):

becuase you move 2x to the other side, then it becomes 16=4x+2x-2 and ya

OpenStudy (anonymous):

me and my fiancé got -6 and we don't know how because if I put that in 2(-6+8) we get 4

OpenStudy (anonymous):

2(-6+8) equals to -12+16 which equals to 4

OpenStudy (anonymous):

@kstmyers 2(x+8) 4x -2 2x + 16 = 4x - 2 2x + 16 - 4x = 4x - 2 - 4x -2x + 16 = -2 -2x + 16 - 16 = -2 - 16 -2x = - 18 -2x/2 = -18/-2 x = 9 <- answer